OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (0, 5) and (–2, 11)?

OpenStudy (anonymous):

y - y1 = m (x - x1)

OpenStudy (anonymous):

get the slope between the two points and use any of those two points to apply in the formula given above.

OpenStudy (anonymous):

the equation of the line that passes through two points (x1,y1) and (x2,y2) is\[\frac{y - y_{1}}{y_{2} - y_{1}}=\frac{x - x_{1}}{x_{2} - x_{1}}\]

OpenStudy (anonymous):

wait